# Stationary solution to Stochastically Forced Euler-Poisson Equations in Bounded Domain: Part 1. 3-D Insulating Boundary

### v2

@@ Abstract.
This paper is concerned with $3$-D stochastic Euler-Poisson equations with insulating boundary conditions forced by the Wiener process. We first establish the global existence and uniqueness of the solution to the system, then we prove that the solution converges to its steady-state time-asymptotically. Compared with the deterministic case, due to the stochastic forces, the momentum does not have the temporal derivative, which causes some difficulty to the energy estimates. To overcome the obstacle caused by the stochastic force, we develop some new techniques to treat the stochastic problems. Furthermore, we observe that the invariant measure is just the Dirac measure generated by the steady-state, in which the time-exponential convergence rate to the steady-state plays an essential role.
@@ 1. Introduction
$\left\{\begin{array}[]{l}\rho_{t}+\nabla\cdot\left(\rho{\bf u}\right)=0,\\ \operatorname{d}_{t}\left(\rho{\bf u}\right)+\left(\nabla\cdot\left(\rho{\bf u}\otimes{\bf u}\right)+\nabla P\left(\rho\right)-\rho\nabla\Phi\right)\operatorname{d}t=-\frac{\rho{\bf u}}{\tau}\operatorname{d}t+\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}_{t}W,\\ \triangle\Phi=\rho-b,\end{array}\right.$ (1.1)
where $\operatorname{d}_{t}$ is the differential notation with respect to time $t$, $\rho$ is the electron density of semiconductors, ${\bf u}$ denotes the particle velocity, $P\left(\rho\right)$ is the pressure, $\Phi$ is the electrostatic potential, $\tau$ is the velocity relaxation time and $b(x)$ is called the doping profile, which is positive and immobile. The above mentioned unknowns $\rho=\rho(\omega,t,x)$, ${\bf u}={\bf u}(\omega,t,x)$, $\Phi=\Phi$, and $P\left(\rho\right)=P\left(\rho(\omega,t,x)\right)$ are stochastic processes as functions with respect to $\omega$, $t$, and $x$, where $\omega$ is a sample in the probability space $\left(\Omega,\mathbb{P}\right)$. For convenience, we use the simplified notions $\rho$, ${\bf u}$, $\Phi$, and $P\left(\rho\right)$ here and hereafter. $W$ is an $\mathcal{H}$-valued cylindrical Brownian motion defined on the filtrated probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, where $\mathcal{H}$ is an auxiliary separable Hilbert space, $\mathcal{F}$ is the filtration, see the definitions of filtration and Wiener process in Appendix 5.
@@
$\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}_{t}W=\sum\limits_{k=1}^{+\infty}\mathbf{F}_{k}\left(\rho,{\bf u}\right)\operatorname{d}_{t}\beta_{k}.$ (1.3)
@@
where ${a_{k}}$ are positive constants, $Y\left(\rho,{\bf u}\right)$ is a smooth function of $\rho$ and ${\bf u}$, satisfying
Here, $\nabla_{\rho,{\bf u}}$ denotes the differential operator with respect to $\rho$ and ${\bf u}$.
@@
which is given in the probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. Here and hereafter, we simply denote the initial data by $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ without confusion.
@@
The hydrodynamic model of semiconductors was first introduced by Blotekjaer [4], which is the deterministically dynamical model presented by Euler-Poisson equations mathematically. For 1-D case, the initial-boundary value problems to Euler-Poisson equations with the insulating boundary and the Ohmic contact boundary were studied by Hsiao-Yang [13], Li-Markowich-Mei [23], respectively, where the solutions are showed to converge to the corresponding subsonic steady-states time-asymptotically, where the doping profile is needed to be flat: $|b^{\prime}(x)|\ll 1$. Such a restriction was then released by Nishibata-Suzuki [27] and Guo-Strass [12] independently. For $N$-D case, Guo-Strass [12] first considered the deterministic 3-D Euler-Poisson equations in bounded domain with insulating boundary, and showed the convergence of solutions to the 3-D subsonic steady-states. Subsequently, Mei-Wu-Zhang [26] investigated the convergence to the steady-states for the n-D radial Euler-Poisson equations with the Ohmic contact boundary. For the whole space without boundary effects, the Cauchy problems to deterministic Euler-Poisson equations were extensively studied in [9, 14, 15, 16, 17, 20]. For the case of free boundary with vacuum, we refer to [24, 25, 29] and the references therein.
@@
When the hydrodynamic model of semiconductors is counted into the stochastic affections, it then becomes the stochastic Euler-Poisson equations with uncertain extra disturbances, see (1.1) with the Wiener process $\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}_{t}W$. This is a new model for semiconductor devices and never touched yet. The main issue of the paper is to investigate this 3-D stochastic Euler-Poisson equations in bounded domain with insulating boundary, and are going to prove the convergence of solutions to the stochastic steady states. The coefficient function of Wiener process $\mathbf{F}_{k}\left(\rho,{\bf u}\right)$, depending on the solutions $\rho$ and ${\bf u}$, is called the multiplicative noise. In most cases, the multiplicative noise magnifies the perturbation and thereby complicating the well-posedness of solutions for evolution systems. The stochastic forces are at most Hölder-$\frac{1}{2}-$continuous in time $t$, resulting in reduced regularity of velocity with respect to time. So from a mathematical standpoint, the study of the stochastic problem helps us to study how the solutions to stochastic Euler-Poisson equations behave in the absence of strong regularity in time. Further, this encourages exploring whether the desirable property remains under the influence of particular types of noise. This is the first attempt to study the asymptotic behavior of solutions to stochastic 3-D Euler-Poisson equations.
For stochastic evolution systems, the solution is called the stationary solution provided that the increment of solutions during evolution is time-independent. Da Prato-Zabczyk [8] listed the basic theory of stationary solutions of general stochastic PDEs in view of invariant measure. As an application of invariant measure for concrete models, Breit-Feireisl-Hofmanová-Maslowski [6] proved the existence of stationary solutions for stochastic compressible Navier-Stokes equations. Gess-Souganidis [11] obtained the large time behavior and existence of invariant measure for the stochastic scalar conservation laws, and the algebraic decay in time holds. In [11], they gave a kind of noise which has better regularization to the problem. Then Dong-Zhang-Zhang [10] proved the existence of stationary solutions to stochastic scalar conservation laws with the multiplicative noise. Bedrossian-Liss [1] gave the existence of stationary measures for stochastic differential equations with a nonlinear term.
@@
such that there hold
@@
Our main result is on the existence of strong solutions near the steady state, and asymptotic stability for insulating boundary condition.
@@
Theorem 1.1.
Let $U$ be a smooth bounded domain in $\mathds{R}^{3}$ and the pressure $P:(0,\infty)\rightarrow(0,\infty)$ be a smooth function, with $P(\cdot)>0$ and $P^{\prime}(\cdot)>0$. Let $\left(\bar{\rho},0,\bar{\Phi}\right)$ be the smooth steady state in Proposition 1.1. There exists a constant $\varepsilon>0$ such that if the initial condition $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ satisfies
$\mathbb{E}\left[\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right]\leqslant\varepsilon^{2m},\quad\forall m\geqslant 2,$ (1.19)
and
@@
then in $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, there exists a unique global-in-time solution $(\rho,{\bf u},\Phi)$ of (1.1):
$\rho,\ {\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right),\ \Phi\in L^{2m}\left(\Omega;C\left([0,T];H^{5}\left(U\right)\right)\right),$ (1.21)
up to a modification. Moreover, there hold the decay rate and the existence of invariant measure:
@@
As mentioned before, the study of stochastic Euler-Poisson equations, totally from the existing studies for the deterministic case, is new and challenging. The idea of the proof is as follows. We first prove the local existence by Banach’s fixed point theorem, then we establish the uniform energy estimates in time $t$ to show the global existence of $\left(\sigma,{\bf u},\phi\right)$. Furthermore, we prove the weighted energy estimates so that we can obtain the asymptotic stability for steady states with the insulating boundary conditions. The a priori estimates imply the tightness of approximates measures, which will converge to an invariant measure by Krylov-Bogoliubov’s theorem in a complete probability space. From the weighted energy estimates, we then prove that the invariant measure for (1.1) is exactly the law of steady state, c.f. Section 4. This intricate relationship has not been uncovered in the asymptotic behavior analysis of stochastic Navier-Stokes equations yet [6].
Here we explain in detail the main difficulties we have to face to and the strategies we are going to propose.
@@
    Weighted energy estimates on account of the estimates of the stochastic integral. Recalling the 3-D deterministic case, for instance, one can multiply the ordinary differential equation of solutions with the exponential function of $t$ directly to facilitate the stability analysis. However, in this paper, in order to estimates the stochastic term, we apply the Burkholder-Davis-Gundy’s inequality to the stochastic integral of the Wiener process. Then the a priori estimates (2.218) is already in the form of time integrals rather than a differential equation. Integration with respect to time twice could not imply the asymptotic stability. Consequently, direct acquisition of asymptotic stability becomes challenging. To overcome this obstacle, we employ the weighted energy estimates. Moreover, the weight is determined by the a priori estimates which should be obtained first, cf. Section 3.1.
@@ 2. Global existence of solutions
In this section, we first establish the local existence of solutions by Banach’s fixed point theorem. Then under the assumption that $\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}$ is small, we get the uniform estimates so as to obtain the global existence of $\rho,{\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right)$.
In form of $\left(\sigma,{\bf u},\phi\right)=\left(\rho-\bar{\rho},{\bf u}-\bar{{\bf u}},\Phi-\bar{\Phi}\right)$, the hydrodynamic system deforms into
$\left\{\begin{array}[]{l}\sigma_{t}+\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)=0,\\ \operatorname{d}_{t}\left({\bf u}\right)+\left(\left({\bf u}\cdot\nabla\right){\bf u}+{\bf u}+\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\operatorname{d}t=\nabla\phi\operatorname{d}t+\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}_{t}W,\\ \triangle\phi=\sigma.\end{array}\right.$ (2.1)
Here we view $\tau$ as a constant without loss of generality for the stability analysis. In terms of component, by Taylor’s expansion, it holds
@@
$\operatorname{d}_{t}\mathbf{w}+\left(\mathcal{A}^{1}\mathbf{w}_{,1}+\mathcal{A}^{2}\mathbf{w}_{,2}+\mathcal{A}^{3}\mathbf{w}_{,3}+\mathcal{B}\mathbf{w}+\mathcal{L}_{{\bf u}}\right)\operatorname{d}t=\mathcal{L}_{\mathcal{\phi}}\operatorname{d}t+f,$ (2.5)
@@
$\mathcal{L}_{{\bf u}}=\left[\begin{array}[]{c}0\\ u^{1}\\ u^{2}\\ u^{3}\end{array}\right],\quad\mathcal{L}_{\mathcal{\phi}}=\left[\begin{array}[]{c}0\\ \triangle^{-1}\sigma_{,1}\\ \triangle^{-1}\sigma_{,2}\\ \triangle^{-1}\sigma_{,3}\end{array}\right],\quad f=-\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}-\mathbb{F}^{1}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,2}-\mathbb{F}^{2}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,3}-\mathbb{F}^{3}\operatorname{d}_{t}W\end{array}\right].$ (2.38)
We define the symmetrizer $\mathcal{D}=\operatorname{diag}\left[Q^{\prime}\left(\bar{\rho}+\sigma\right),\bar{\rho}+\sigma,\bar{\rho}+\sigma,\bar{\rho}+\sigma\right]:=\operatorname{diag}\left[d_{1},d_{2},d_{3},d_{4}\right]$, then the system deforms into
$\mathcal{D}\operatorname{d}_{t}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}+\tilde{\mathcal{L}}_{{\bf u}}\right)\operatorname{d}t=\tilde{\mathcal{L}}_{\phi}\operatorname{d}t+\tilde{f},$ (2.39)
@@
$\tilde{f}=-\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}\operatorname{d}t-\mathbb{F}^{1}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,2}\operatorname{d}t-\mathbb{F}^{2}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,3}\operatorname{d}t-\mathbb{F}^{3}\operatorname{d}_{t}W\end{array}\right].$ (2.73)

### v3

@@ Abstract.
This paper is concerned with $3$-D stochastic Euler-Poisson equations with insulating boundary conditions forced by the Wiener process. We first establish the global existence and uniqueness of the solution to the system, then we prove that the solution converges to its steady-state time-asymptotically. To obtain the converging rate, we need to develop weighted energy estimates, which are not required for the deterministic counterpart of the problem. Moreover, we observe that the invariant measure is just the Dirac measure generated by the steady-state, in which the time-exponential convergence rate to the steady-state plays an essential role.
@@ 1. Introduction
$\left\{\begin{array}[]{l}\rho_{t}+\nabla\cdot\left(\rho{\bf u}\right)=0,\\ \operatorname{d}\left(\rho{\bf u}\right)+\left(\nabla\cdot\left(\rho{\bf u}\otimes{\bf u}\right)+\nabla P\left(\rho\right)-\rho\nabla\Phi\right)\operatorname{d}t=-\frac{\rho{\bf u}}{\tau}\operatorname{d}t+\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}W,\\ \triangle\Phi=\rho-b,\end{array}\right.$ (1.1)
where ``$\operatorname{d}"$ in (1.1) is the differential notation with respect to time $t$, in comparison to gradient $\nabla$ and Laplacian $\triangle$ for spatial derivatives, $\rho$ is the electron density of semiconductors, ${\bf u}$ denotes the particle velocity. $P\left(\rho\right)$ is the pressure, $\Phi$ is the electrostatic potential, $\tau$ is the velocity relaxation time and $b(x)$ is called the doping profile, which is positive and immobile. The above mentioned unknowns $\rho=\rho(\omega,t,x)$, ${\bf u}={\bf u}(\omega,t,x)$, $\Phi=\Phi(\omega,t,x)$, and $P\left(\rho\right)=P\left(\rho(\omega,t,x)\right)$ are stochastic processes as functions with respect to $\omega$, $t$, and $x$, where $\omega$ is a sample in the complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. For convenience, we use the simplified notions $\rho$, ${\bf u}$, $\Phi$, and $P\left(\rho\right)$ here and hereafter. $W$ is an $\mathcal{H}$-valued cylindrical Brownian motion defined on the filtrated probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, where $\mathcal{H}$ is an auxiliary separable Hilbert space, $\mathcal{F}$ is the filtration, see the definitions of filtration and Wiener process in Appendix 5.
@@
$\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}W=\sum\limits_{k=1}^{+\infty}\mathbf{F}_{k}\left(\rho,{\bf u}\right)\operatorname{d}\beta_{k}e_{k}.$ (1.3)
@@
where ${a_{k}}$ are positive constants, $Y\left(\rho,{\bf u}\right)$ is a smooth function of $\rho$ and ${\bf u}$, and can be bounded by the homogeneous polynomials.
@@
which is given in the probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, $\rho_{0}\left(\omega,x\right)>0$. Here and hereafter, we simply denote the initial data by $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ without confusion.
@@
The hydrodynamic model of semiconductors was first introduced by Blotekjaer [4], which is the deterministically dynamical model presented by Euler-Poisson equations mathematically. For 1-D case, the initial-boundary value problems to Euler-Poisson equations with the insulating boundary and the Ohmic contact boundary were studied by Hsiao-Yang [29], Li-Markowich-Mei [39], respectively, where the solutions are showed to converge to the corresponding subsonic steady-states time-asymptotically, where the doping profile is needed to be flat: $|b^{\prime}(x)|\ll 1$. Such a restriction was then released by Nishibata-Suzuki [44] and Guo-Strauss [21] independently. For $N$-D case, Guo-Strauss [21] first considered the deterministic 3-D Euler-Poisson equations in bounded domain with insulating boundary, and showed the convergence of solutions to the 3-D subsonic steady-states. Subsequently, Mei-Wu-Zhang [42] investigated the convergence to the steady-states for the $N$-D radial Euler-Poisson equations with the Ohmic contact boundary. For the whole space without boundary effects, the Cauchy problems to deterministic Euler-Poisson equations were extensively studied in [10, 30, 31, 32, 33, 36]. For the case of free boundary with vacuum, we refer to [40, 43, 51] and the references therein. For the formulation of singularities in compressible Euler-Poisson equations and the large time behavior of Euler equations with damping, one can refer to [49] and [50], respectively.
@@
When the hydrodynamic model of semiconductors is counted into the stochastic affections, it then becomes the stochastic Euler-Poisson equations with uncertain extra disturbances, see (1.1) with the Wiener process $\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}W$. This is a new model for semiconductor devices and never touched yet. The main issue of the paper is to investigate this 3-D SEP in bounded domain with insulating boundary, and are going to prove the convergence of solutions to the stochastic steady states. The coefficient function of Wiener process $\mathbf{F}_{k}\left(\rho,{\bf u}\right)$, depending on the solutions $\rho$ and ${\bf u}$, is called the multiplicative noise. In most cases, the multiplicative noise magnifies the perturbation and thereby complicating the well-posedness of solutions for evolution systems. The stochastic forces are at most Hölder-$\frac{1}{2}-$continuous in time $t$, resulting in reduced regularity of velocity with respect to time. So from a mathematical standpoint, the study of the stochastic problem helps us to study how the solutions to stochastic Euler-Poisson equations behave in the absence of strong regularity in time. Further, this encourages exploring whether the desirable property remains under the influence of particular types of noise. This is the first attempt to study the asymptotic behavior of solutions to stochastic 3-D Euler-Poisson equations.
For stochastic evolution systems, the solution is called the stationary solution provided that the increment of solutions during evolution is time-independent. Originally, the study of stationary measures dates back to the works of Hopf [28], Doeblin [9], Doob [12], Halmos [22, 23], Feller [13], and Harris and Robbins [25, 24], who contributed to the theory of discrete Markov processes from 1930s to 1950s. The study of invariant measure of fluid models dates back to Cruzeiro [8] for stochastic incompressible Navier-Stokes equations in 1989, by Galerkin approximation with dimensions $D\geqslant 2$. Flandoli [14] proved existence of an invariant measure by the “remote start” method for 2-D incompressible Navier-Stokes equations in 1994. One year later Flandoli-Gatarek [15] showed existence of stationary solution for 3-D incompressible Navier-Stokes equations by a different method with [8]. In 2002, Mattingly [41] proved the existence of exponentially attracting invariant measure with respect to initial data, for incompressible N-S equations. Later, Goldys-Maslowski [20] showed that transition measures of the 2-D stochastic Navier-Stokes equations converge exponentially fast to the corresponding invariant measures in the distance of total variation. Then for 3-D case, Da Prato and Debussche [46] constructed a transition semigroup for 3-D stochastic Navier-Stokes equations without the uniqueness, which allows for rather irregular solutions. Flandoli-Romito [17] used the classical Stroock-Varadhan type argument to find the almost sure Markov selection. The above works are for the incompressible case. For stochastic compressible Navier-Stokes equations, Breit-Feireisl-Hofmanová-Maslowski [6] proved the existence of stationary solutions. Compared to Navier–Stokes equations, the regularity effect of viscosity is lost for Euler system. Hofmanová-Zhu-Zhu [27] selected the dissipative global martingale solutions to the stochastic incompressible Euler system, and obtained the non-uniqueness of strong Markov solutions. Very recently, they [26] showed that stationary solution to the Euler equations is a vanishing viscosities limit in law of stationary analytically weak solutions to Navier-Stokes equations. In terms of the non-uniqueness studies, some scholars believe that a certain stochastic perturbation can provide a regularizing effect of the underlying PDE dynamics. For instance, Flandoli-Luo [16] showed that a noise of transport type prevents a vorticity blow-up in the incompressible Navier-Stokes equations. A linear multiplicative noise prevents the blow up of the velocity with high probability for the 3-D Euler system, which was shown by Glatt-Holtz-Vicol [19]. Gess-Souganidis [18] investigated the large-time behavior and established the existence of an invariant measure for stochastic scalar conservation laws, demonstrating that an algebraic decay rate in time holds. In their work, they introduced a particular type of noise that provided stronger regularization properties for the problem. Then Dong-Zhang-Zhang [11] proved the existence of stationary solutions with the multiplicative noise. For stochastic conservation laws, Da Prato-Gatarek studied the existence and uniqueness of invariant measure for stochastic Burgers equation [47]. Da Prato-Zabczyk listed the basic theory of stationary solutions of general stochastic PDEs in view of invariant measure in book [48]. Bedrossian-Liss [1] gave the existence of stationary measures for stochastic ordinary differential equations with a nonlinear term. To the best of our knowledge, the stationary solutions of SEP have not been explored previously. For our SEP, the electrostatic potential term $\rho\nabla\Phi\operatorname{d}t$ and the relaxation term $\frac{\rho{\bf u}}{\tau}\operatorname{d}t$ are actually damping terms providing better regularity than Euler equations. In this paper, we could show the existence and uniqueness of invariant measure in more regular space.
@@
such that there holds
@@
Our main result is on the existence of solutions near the steady state, and asymptotic stability for insulating boundary condition.
@@
Theorem 1.1 (Global existence).
Let $U$ be a smooth bounded domain in $\mathds{R}^{3}$ and the pressure $P:(0,\infty)\rightarrow(0,\infty)$ be a smooth function, with $P(\cdot)>0$ and $P^{\prime}(\cdot)>0$. Let $\left(\bar{\rho},0,\bar{\Phi}\right)$ be the smooth steady state in Proposition 1.1. and
@@
then in $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, there exists a unique global-in-time strong solution $(\rho,{\bf u},\Phi)$ to the initial and boundary problem (1.1)-(1.5)-(1.6):
$\rho,~{}{\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right),\quad\Phi\in L^{2m}\left(\Omega;C\left([0,T];H^{5}\left(U\right)\right)\right),\forall~{}T>0,$ (1.19)
up to a modification, for any fixed integer $m\geqslant 2$.
Moreover, for the small perturbation problem, there hold the the existence of invariant measure and decay rate.
Theorem 1.2 (Convergence to steady state).
Assume that the stochastic forces satisfies
Here, $\nabla_{\rho,{\bf u}}$ denotes the differential operator with respect to $\rho$ and ${\bf u}$. If there exists a constant $\varepsilon>0$ such that the initial condition $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ satisfies (1.18) and
$\mathbb{E}\left[\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right]\leqslant\varepsilon^{2m},\quad\forall~{}m\geqslant 2,$ (1.21)
then there hold the decay rate and the existence of invariant measure:
@@
As mentioned before, the study of stochastic Euler-Poisson equations, totally from the existing studies for the deterministic case, is new and challenging. The idea of the proof is as follows. We first prove the local existence by Banach’s fixed point theorem, then we establish the uniform energy estimates in time $t$ to show the global existence of $\left(\sigma,{\bf u},\phi\right)$. Furthermore, we prove the weighted energy estimates so that we can obtain the asymptotic stability for steady states with the insulating boundary conditions. The a priori estimates imply the tightness of approximates measures, which will converge to an invariant measure by Krylov-Bogoliubov’s theorem in a complete probability space. The global existence does not require the small perturbation condition (1.20) and (1.21). However, the existence of invariant measure in Theorem 4.2, requires (1.20) and (1.21). From the weighted energy estimates, we then prove that the invariant measure for (1.1) is exactly the law of steady state, c.f. Section 4. This intricate relationship has not been uncovered in the asymptotic behavior analysis of stochastic Navier-Stokes equations yet [41, 20].
Here we explain in detail the main difficulties we face to and the strategies we are going to propose.
@@
    Weighted energy estimates on account of the estimates of the stochastic integral. Recalling the 3-D deterministic case [21], for instance, based on the energy estimates, one can obtain ordinary differential inequality (ODI). Then they multiply the ODI with the exponential function of $t$ directly to facilitate the stability analysis. However, in this paper, in order to estimates the stochastic term, we apply the Burkholder-Davis-Gundy’s inequality to the stochastic integral of the Wiener process. Then the a priori estimates (2.221) is already in the form of time integrals rather than an ODI. Integration with respect to time twice could not imply the asymptotic stability. Consequently, direct acquisition of asymptotic stability becomes challenging. To overcome this obstacle, we employ the weighted energy estimates. Moreover, the weight is determined by the a priori estimates which should be obtained first, cf. Section 3.1.
@@ 2. Global existence of solutions
In this section, we first establish the local existence of strong solutions by Banach’s fixed point theorem. Specifically, we derive the system of perturbed solutions in matrix notation by (1.1). In Step 1, We symmetrize it to simplify the energy estimates and proceed to linearize the system. In Step 2, following a standard procedure in view of Picard interation, we establish the uniform estimates onto mapping. We utilize Itô’s formula and the Burkholder-Davis-Gundy inequality to estimate the stochastic force. In Step 3, we demonstrate contraction. In Step 4, we get the a priori estimates in §2.2 so as to obtain the global existence of $(\rho-\bar{\rho},{\bf u})$, or equivalently, $\rho,{\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right)$ in §2.3. Step 5 is about the proof of global existence.
In form of $\left(\sigma,{\bf u},\phi\right)=\left(\rho-\bar{\rho},{\bf u},\Phi-\bar{\Phi}\right)$, the hydrodynamic system deforms into
$\left\{\begin{array}[]{l}\sigma_{t}+\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)=0,\\ \operatorname{d}\left({\bf u}\right)+\left(\left({\bf u}\cdot\nabla\right){\bf u}+{\bf u}+\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\operatorname{d}t=\nabla\phi\operatorname{d}t+\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}W,\\ \triangle\phi=\sigma.\end{array}\right.$ (2.1)
Here we view $\tau$ as a constant $1$ without loss of generality for the stability analysis. In terms of component, by Taylor’s expansion, it holds
@@
$\operatorname{d}\mathbf{w}+\left(\mathcal{A}^{1}\mathbf{w}_{,1}+\mathcal{A}^{2}\mathbf{w}_{,2}+\mathcal{A}^{3}\mathbf{w}_{,3}+\mathcal{B}\mathbf{w}+\mathcal{L}_{{\bf u}}\right)\operatorname{d}t=\mathcal{L}_{\mathcal{\phi}}\operatorname{d}t+f,$ (2.5)
@@
$\mathcal{L}_{{\bf u}}=\left[\begin{array}[]{c}0\\ u^{1}\\ u^{2}\\ u^{3}\end{array}\right],\quad\mathcal{L}_{\mathcal{\phi}}=\left[\begin{array}[]{c}0\\ \triangle^{-1}\sigma_{,1}\\ \triangle^{-1}\sigma_{,2}\\ \triangle^{-1}\sigma_{,3}\end{array}\right],\quad f=-\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}-\mathbb{F}^{1}\operatorname{d}W\\ h\left(\sigma\right)_{,2}-\mathbb{F}^{2}\operatorname{d}W\\ h\left(\sigma\right)_{,3}-\mathbb{F}^{3}\operatorname{d}W\end{array}\right].$ (2.38)
Step 1: Symmetrizing and Linearizing.
We define the symmetrizer $\mathcal{D}=\operatorname{diag}\left[Q^{\prime}\left(\bar{\rho}+\sigma\right),\bar{\rho}+\sigma,\bar{\rho}+\sigma,\bar{\rho}+\sigma\right]:=\operatorname{diag}\left[d_{1},d_{2},d_{3},d_{4}\right]$. Then the system deforms into
$\mathcal{D}\operatorname{d}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}+\tilde{\mathcal{L}}_{{\bf u}}\right)\operatorname{d}t=\tilde{\mathcal{L}}_{\phi}\operatorname{d}t+\tilde{f},$ (2.39)
@@
$\tilde{f}=-\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}\operatorname{d}t-\mathbb{F}^{1}\operatorname{d}W\\ h\left(\sigma\right)_{,2}\operatorname{d}t-\mathbb{F}^{2}\operatorname{d}W\\ h\left(\sigma\right)_{,3}\operatorname{d}t-\mathbb{F}^{3}\operatorname{d}W\end{array}\right].$ (2.73)